The graph of the function f(x)=mx+b contains the points (r,3) and (7,s). Express s in terms of r if the graph is parallel to the line 3x-2y=7. The answer is s=-3/2r+27/2 I just don't know how to get there thank you
Answer by Alan3354(69443)

The graph of the function f(x)=mx+b contains the points (r,3) and (7,s). Express s in terms of r if the graph is parallel to the line 3x-2y=7. The answer is s=-3/2r+27/2
----------
Parallel lines have the same slope, m.
--
Find the slope of the line 3x-2y=7
Put it in slope-intercept form, y = mx + b
That means solve for y.
3x-2y=7
3x = 2y + 7
2y = 3x -7
y = (3/2)x - 7/2
Slope = 3/2
----------
Find the slope of the 2 points.
Slope = diffy/diffx = (s-3)/(7-r)
So (s-3)/(7-r) = 3/2
Cross multiply.
2(s-3) = 3(7-r)
2s-6 = 21-3r
2s = -3r + 27
s = (-3/2)r + 27/2